OpenStudy (anonymous):

Write the formula given the following arithmetic sequences. 1, -1, -3, -5, ... Select one: a. tn = 1 + 2(n – 1) b. tn = 1 + (n – 2) c. tn = 1 - 2(n – 1) d. tn = 2 - (n – 1)

OpenStudy (anonymous):

do you get the common difference?

OpenStudy (anonymous):

c.d = t(n+1)-t(n) => -1-1 = -3 - (-1) = -5-(-3) ..

OpenStudy (anonymous):

omg. what?

OpenStudy (anonymous):

Oh, Do you know know what the common difference is in an arithmetic progression?

OpenStudy (anonymous):

no i dont,

OpenStudy (anonymous):

Okay.. So, An arithmetic progression is called so because it forms a series such that there is a definite difference between two consecutive numbers .. In the series that you gave above, 1, -1, -3, -5, ... The common difference is : (-1-1) = (-3+1) = (-5+3 .... and so one.. Therefore, the common difference is -2. from there on, just apply the formula as : tn = a+(n-1)d where, n is the number of terms and d is the common difference..

OpenStudy (anonymous):

so i understand it better but i still don't see the answer. you say d is -2 but when i look up to the answers there is no -2 in that position. so im still confused i guess. thanks for helping thou.

OpenStudy (anonymous):

a. tn = 1 + 2(n – 1) b. tn = 1 + (n – 2) c. tn = 1 - 2(n – 1) d. tn = 2 - (n – 1) Those are your options. Now we get d as -2, agreed? and the equation (gen) is - a+d(n-1) {your series was - 1,-1,-3,-5} .. therefore a=1 (first term) Compare the gen eq and the options : so, tn = 1 + (-2)(n-1) = 1 - 2(n-1) , Got it now? Any confusion? Feel free to ask :)

OpenStudy (anonymous):

cool thanks.(: i undertand now,